One spring I placed inside the other and compressed between a pair of parallel plates until the distance between the two plates is 40mm calculate the force applied to the spring assembly



1
Expert's answer
2021-07-12T03:36:51-0400

Force applied to the spring assembly


"F =\frac{1}{2}QE"


V = 0.8


"Q = CV =1 \times 0.8 = 0.8"


"E = \frac{V}{d} =\frac{0.8}{0.04}= 20"


"F =\frac{1}{2}QE = \frac {1}{2}\times20\times0.8 = 8N"